OpenStudy (anonymous):

find all points of absolute minima and maxima on the given closed interval. y = -x^2 + 4x - 3; [0,3]

OpenStudy (anonymous):

we can draw the graph first 1) find x-intercept. (put y=0) put y=0, -x^2 +4x -3=0 x^2 -4x+3=0 (x-1)(x-3)=0 x=1 or 3 2) find y-intercept (put x=0) put x=0, y=-0+0-3 y=-3 3) Find the vertex of the curve (by completing square) y=-x^2+ 4x -3=-(x^2 -4x)-3=-(x^2-4x+4)+4-3=-(x-2)^2 +1 so the vertex is (2,1) 4) find the opening of the curve. as the coefficient of x^2 is -ve, the curve open downward. Now, we can draw the curve.  For the range given, [0,3] for [], it means that the end points are also included. Can you now find the max and min from the graph? If yes, give me the answer. I can tell you if it is right or wrong. If no, tell me what's your concern.